Sector 7 Useful Operational Data for Run 2 of FY06
written by Sector 7 staff.
Started Jan 30, 2005
Table of Content:
The work shown below is an account of useful operational data taken during
run 2 of FY06 on 7ID also called 2006-1 by APS. This page will contain useful
stability information on the 7ID High Heat Load Monochromator and also on the
various repairs and problems identified during the run. It first details the
Summary of the Jan 2006 shutdown activities.
Pictures from the January 2006 shutdown.
This shutdown, the monochromator was warmed up and the new cryocooler lines
were installed through the new labyrinth (see pictures soon).
Bernhard Adams and ED installed new Compton shields inside the 7ID-A High Heat
load monochromator. One of these is now on top of the first crystal. The second
one shields all the mechanics of the second crystal assembly(see pictures
above). A more detailed account is shown in an Elog here.
I also remembered recently that one can use the beamline monochromator on Si
(333). This has not been used or tested by the beamline staff but it may lead
to improved energy resolution. A more detailed account is shown in an Elog here.
Jan 31, 2006.
The beamline started to operate today at 11h24am. The start was delayed due to
the fact that several beamlines were ahead of 7ID for their type-C surveys.
The mono was also misaligned. So it wasn't until 2pm that we had beam in the C
hutch. The mini-hutch after ID-B was opened during the shutdown to put new RED
configuration control tag inside. So a type C was completed in mono mode.
Useful ELOGs from the winter 2005 run (local access only)
Lessons learned during run 2 of FY06